OpenStudy (anonymous):

8th grade math (easy) Medal and Fan!!! /just have to find where the first error is/ ..Systems of Equations..

OpenStudy (anonymous):

Equation 1: m = 8 + 2n Equation 2: 4m = 4 + 4n Step 1: –4(m) = –4(8 + 2n) [Equation 1 is multiplied by –4.] 4m = 4 + 4n [Equation 2] Step 2: –4m = –32 – 8n [Equation 1 in Step 1 is simplified.] 4m = 4 + 4n [Equation 2] Step 3: –4m + 4m = –32 – 8n + 4n [Equations in Step 2 are added.] Step 4: 0 = –32 – 4n Step 5: n = –8 In which step did the student first make an error? (1 point) Step 1 Step 2 Step 3 Step 4

OpenStudy (welshfella):

a number is left out in the add - Can you see which number it is?

OpenStudy (anonymous):

2?

OpenStudy (welshfella):

No - there is no invilved

OpenStudy (welshfella):

* no 2 involved

OpenStudy (anonymous):

I don't believe I understand then... are you saying there is a number missing in Step 3?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

4!

OpenStudy (welshfella):

thats correct

OpenStudy (anonymous):

ohh, ok so the answer is Step 3?

OpenStudy (welshfella):

yes it should be -4m + 4m = -32 - 8n + 4 + 4n

OpenStudy (anonymous):

oh ok, thanks for explaining. I like when people explain instead of just giving me the answer. Im ur fan now, and I gave u a medal!

OpenStudy (welshfella):

ty yw Yes giving the answer doesn't teach you much.